OpenStudy (anonymous):

The digits 0, 1, 2, 3, and 4 are to be used to form a five-digit number. How many different numbers can be formed if a) repetitions are permitted? b) repetition is not allowed? c) the number must be odd and repetitions are not allowed? d) the number must be greater than 30,000 and repetition is not allowed?

OpenStudy (anonymous):

Can you please help me @iGreen Thank you :)

OpenStudy (dan815):

okay repetitions are allowed _ _ _ _ _ for the first digit we can pick 1 to 4 the 2nd to last digit we can pick 0 to 4 so 4*5*5*5*5 = total 5 digit numbers with repetition allowed

OpenStudy (dan815):

repetitions not allowed _ _ _ _ _ we still have 4 numbers to pick from 1 to 4 in the first part now we have 3 of those numbers left and the 0 number so we have 4 choices to pick from in the 2nd one and then 3 and the 2 and 1 4*4*3*2*1= total number of 5 digit numbers with no repetition is allowed

OpenStudy (dan815):

number must be odd so the last digit has to end with 1 or 3 only 2 choices for the last digit, then we have 4 choices left for the 2nd last then 3, then 2 then 1 _ _ _ _ _ 1*2*3*4*2 = total number of odd digits with no repetition

OpenStudy (dan815):

for a number to be greater than 30000 this means the first digit is 3 or 4 so 2 choices and no repeitions are allowed so _ _ _ _ _ 2*4*3*2*1= total number of digits greater than 30000, with no repetition allowed

OpenStudy (dan815):

Do you understand >_>

OpenStudy (dan815):

you've been awfully quiet

OpenStudy (anonymous):

@dan815 thank you so much :D sorry my internet connection is very slow i had to refresh it mulitple times :))

OpenStudy (anonymous):

in a). why is that only 1-4 can be used in the first digit?

OpenStudy (kainui):

If you had 0 for the first digit it wouldn't be a 5-digit number, it'd be a 4-digit number!